Hi, can someone please help me solve the following...
x^2 + y^2 = 14,
x^2 - y^2 = 4

Answer by stanbon(75887)

x^2 + y^2 = 14,
x^2 - y^2 = 4
--------------
Add and solve for "x^2":
2x^2 = 18
x^2 = 9
x = +3 or x = -3
---
If x = +3 or -3 solve for "y":
y^2 = 14-9
y^2 = 5
y = sqrt(5) or y = -sqrt(9)
----
Ans: (3,sqrt(5))or (-3,sqrt(5)) or (3,-sqrt(5)) or (-3,-sqrt(5))
=====================
Cheers,
Stan H.
=====================